# MFTF: Mask-free Training-free Object Level Layout Control Diffusion Model

### v1

@@ 4.1 Dynamic Cross-attention Mask Generation
The threshold $\eta$ has a significant impact on the generated images through the attention mask. When $\eta=0$, no attention mask is applied to the query, making it impossible to separate the object from the background. When $\eta=1$, the query is entirely removed from the self-attention process. For multiple-object position control, the threshold $\eta$ controls the degree to which the original object is retained.
The cross-attention maps corresponding to individual tokens $i$ are illustrated in Fig. 4, which encapsulates both semantic and structural information. However, the structural information corresponding to each text token overlaps significantly, which complicates achieving precise absolute layout control. For instance, the cross-attention map in Fig. 4 generated for the prompt “a small cat sitting on a desk”, the token “cat” captures the majority of the structural information for the cat object. However, other tokens such as “small”, “sitting”, “on”, and “desk” also contribute overlapping profile information related to the “cat”, thereby demonstrating the inherent redundancy in the attention map.
Figure 3: The cross-attention mask $M^{i}_{s}$ is generated from the cross-attention layers $A^{i}_{s}$
@@ 4.2 Self-Attention-Based Layout Control
The self-attention layer $Q$ encodes both detailed object information and global layout characteristics of the generated images. The semantic and structural information represented by $Q$ varies across different self-attention layers. Specifically, $Q_{s,l}$ corresponds to the self-attention representation at layer $l$ as illustrated in Fig. 5. Lower layers ($l=0$ to $5$) lack significant semantic information but begin to establish the foundational structure. In the intermediate layers ($l=6$ to $10$), structural features become more prominent and start to take shape. The upper layers ($l=11$ to $15$) contain detailed object-specific structural and semantic information. For example, as shown in Fig. 6, layer $l=10$ encapsulates information connecting the object to its background, layer $l=13$ focuses on the object’s overall profile, and layer $l=15$ captures fine details of the generated image.
Figure 5: The visualization of $Q$ is presented at different self-attention layers $l=5,10,15$, showing the effects of the following conditions: without the application of the cross-attention mask, with the cross-attention mask applied, and with the additional application of positional control $L$. Furthermore, the residual background in $Q$ after the removal of objects is also displayed.
Figure 6: The visualization of $Q$ is presented at different self-attention layers $l=11,13,15$, showing the effects of the following conditions: without the application of the cross-attention mask, with the cross-attention mask applied, and with the additional application of positional control $L$. Furthermore, the residual background in $Q$ after the removal of objects is also displayed.
The cross-attention mask effectively selects objects from the background following Eq. 4, and is shown in the second column of Fig. 5 and Fig. 6.
@@
The layout of these objects is then controlled using the layout control parameters $L$ using Eq. 5, allowing the manipulation of object locations in $Q_{s}$. This process generates the modified query $Q_{ms}$, as illustrated in the third column of Fig. 5 and Fig. 6.
@@
In contrast to the multi-self-attention mechanism in Masactrl [2], which utilizes $K_{s}$ and $V_{s}$ for the source diffusion model, the MFTF model replaces $Q_{t}$ with $Q_{ms}$ in the target diffusion model to calculate self-attention, as illustrated in Fig. 7.
Figure 7: Modified multi-self-attention.
@@ 4.3.1 Layout Control and Semantic Editing
The model successfully demonstrates concurrent single-object positional control and semantic editing, as illustrated in Fig. 8. The first column presents images generated from the source prompts: ”a small cat sitting on a desk” ”a cat sitting on a chair, no background” and ”a painting on the wall in a cozy house”. Following the application of layout controls $L$, the resulting images are shown in the second, third, and fourth columns. These results highlight the ability to apply positional control, including translation, rotation, and scaling, both independently and simultaneously for target image generation. Object-level layout control and semantic editing are applied together, with modified target prompts such as ”a small cat sitting on grass” ”a small cat sitting on grass” and ”a painting on a chair”. The images generated from these prompts are displayed in the last column of Fig. 8.
Figure 8: MFTF manipulation of the object’s spatial positioning (translation, rotation, scaling) while simultaneously enabling semantic modifications.
@@ 4.3.2 Multiple Object Layout Control
The model is capable of controlling the positions of multiple objects simultaneously, as demonstrated in Fig. 9. The first image in the first row, generated from the source prompt $P_{s}$ = ”a small car and a cabin in the style of Disney, artstation”, shows the initial configuration. In this case, the ”car” on the left-hand side is translated to the left, while the ”cabin” is translated to the right. Subsequently, the ”car” is rotated by $30^{\circ}$ and $330^{\circ}$ clockwise. The second to fourth images in the first row show the target images after the application of layout control $L$. The corresponding $Q_{s}$ and $Q_{ms}$ values are presented in the second row. The third row illustrates another example of concurrent object position control for multiple objects.
@@
Figure 9: MFTF multi-object layout control. Demonstration of concurrently control the positions and arrangements of multiple objects within a scene, enabling independent adjustments to the layout of each object while preserving overall structural coherence.
4.3.3 Text-Controlled Semantic Segmentation
The attention map $M_{s}$ and the masked queries, $Q_{s}$ or $Q_{\text{background}}$, can also be utilized for text-guided image segmentation, as illustrated in Fig. 10. Here, $M_{s}$ serves as the mask for the token ”panda”, which emphasizes the object’s profile, while $Q_{s}$ or $Q_{\text{background}}$ is used to generate detailed feature segmentation.
@@ 4.3.3 Text-Controlled Semantic Segmentation
Figure 10: The MFTF model enables text-controlled image segmentation through the generation of segmentation masks. Segmentation cross-attention mask $M_{s}$ and the masked query $Q_{ms}$ for token “panda” and “bird”.
4.4 Ablation Study
The procedure for applying layout control includes the initial steps up to step $t^{*}$, after which these steps are omitted until the final step. As shown in Fig. 11, the initial steps determine the positioning and arrangement of objects within the structure, while the later denoising steps focus on generating structural details. Layout control is effective when the initial denoising steps are included. However, if these initial steps are omitted, layout control fails to operate as intended.
@@ 4.4 Ablation Study
The initial layers do not contain meaningful structural information; rather, detailed structural information gradually emerges in the higher layers of the U-Net architecture, as shown in Fig. 12. Applying the cross-attention mask to all self-attention layers yields the best position control. However, applying the mask only to the higher-level layers also achieves effective control. Conversely, when the attention mask is applied to a smaller number of self-attention layers, the effectiveness of layout control diminishes, ultimately resulting in a complete loss of positional control.
@@
4.5 Limitations
Despite the effectiveness of the proposed model in achieving single-object and multi-object layout control, as well as simultaneous layout control and semantic editing, several limitations remain. A primary challenge arises from the mixed structural information present in the cross-attention layer, which makes it difficult to fully encapsulate complete structural information within a single attention mask. This limitation hinders the model’s ability to fully separate objects from the background. Additionally, in multi-object position control, the model struggles with separate objects with complex interrelationships, particularly in scenarios involving tightly coupled or overlapping objects.
5 Conclusion
In this paper, we have presented a mask-free, fine-tuning-free object-level layout control diffusion model. The model utilizes self-generated attention masks and applies cross-attention masks to effectively separate objects from the background within the self-attention layers. By leveraging multi-cross attention, the model controls the layout of objects, enabling both single-object and multi-object layout control. MFTF supports concurrent layout control and semantic editing, as well as text-guided image segmentation, demonstrating its promising potential for further research and practical applications in this domain. ( to be modified) Future applications of MFTF models could potentially aid humans in creating compelling custom images with unprecedented speed and ease.

### v2

@@ 4.1 Cross-Attention Mask Creation
The model dynamically generates cross-attention masks $M^{i}_{s}$ for token $i$ at each sampling step $t$ by leveraging the definition of the cross-attention map. The cross-attention map, represented as $A^{i}_{s}=QK^{T}$, has dimensions $b,n,m$, where $b$ corresponds to the batch size, $n$ is the feature sequence length, and $m$ is the text sequence length corresponds to token $0-m$. This enables controlled image generation by updating the cross-attention masks during each sampling step, ensuring layout control throughout the generation process.
@@ 4.2 Self-Attention-Based Layout Control
The self-attention layer $Q$ encodes both detailed object information and global layout characteristics of the generated images. The semantic and structural information represented by $Q$ varies across different self-attention layers. Lower layers ($l=0$ to $5$) lack significant semantic information but begin to establish the foundational structure. In the intermediate layers ($l=6$ to $10$), structural features become more prominent and start to take shape. The upper layers ($l=11$ to $15$) contain detailed object-specific structural and semantic information. $Q_{s,l}$ corresponds to the self-attention layer $l$ as illustrated in SFig. 10.
The cross-attention mask enables effective objects selection from the background following Eq. 4, and is shown in the second column of Fig. 4.
@@
The layout of these objects is then controlled using the layout control parameters $L$ following Eq. 5, allowing the manipulation of object locations in $Q_{s}$. This process generates the modified query $Q_{ms}$, as illustrated in the third column of Fig. 4.
@@
In contrast to the multi-self-attention mechanism in Masactrl [2], which utilizes $K_{s}$ and $V_{s}$ for the source diffusion model, MFTF model replaces $Q_{t}$ with $Q_{ms}$ in the target diffusion model to calculate self-attention, as illustrated in SFig. 8.
4.3 Experiments
4.3.1 Experimental Setup
We implement the proposed method using the state-of-the-art Stable Diffusion text-to-image model [15], leveraging the publicly available v1.4 checkpoints. Sampling is performed using DDIM [19] with $30$ steps in total, and the classifier guidance scale set to $7.5$. The Gaussian random noise is initialized as $Z^{T}_{s}=Z^{T}_{t}$ for the source and target diffusion models Layout control is applied during steps $0\sim 15,20$ across layers $0\sim 16$. The threshold for mask generation ranges from $0.1$ to $0.3$ or is set to $1$ for objects to be dropped.
Figure 6: MFTF object spatial positioning (translation, rotation, scaling) and simultaneously semantic modifications.
4.3.2 Qualitative Evaluation
The quantitative comparison of images generated using text descriptions alone for object-level layout control and those generated by MFTF model presented in the first and second rows of Fig. 5. The source images relies solely on text descriptions fails to control object positions accurately. In contrast, MFTF model successfully positions objects by incorporating additional layout control parameters. For example, in the first pair of images, the generated image incorrectly places the “plant” on the left side instead of the intended right side. The plant is correctly positioned on the right in MFTF model. Similarly, in the last pair of images in the second row, the generated image places “strawberry” inside the mug on the desk instead of on the desk itself. With MFTF model, the ”strawberry” is accurately positioned on the desk as intended. The quantitative comparison of images generated with and without MFTF model for layout control is shown in the third and fourth rows of Fig. 5. In this case, the text description does not provide specific positional control, but users may still want to adjust the positions of objects to better align with their intended layout. For example, the “bus” is repositioned deeper into the forest, the “moon” shrinks in size, the “scientist” is rotated by 30 degrees, and the “fairplace” is shown from a front view.
4.3.3 Quantitative Evaluation
Quantitative comparison of 30 image pairs generated by the Stable Diffusion model and the MFTF model is presented in Tab. 1. In this table, the LPIPS score measures the feature-level similarity between images generated by the two models. The results indicate that feature-level similarity is well-preserved after applying layout control. Additionally, the CLIP score, also shown in Tab. 1, evaluates the text-image alignment for both models. The minimal variation in the CLIP score demonstrates that the layout control process has negligible impact on text-image alignment, ensuring and improving consistency in semantic interpretation after layout adjustments.
Model CLIP Score LPIPS
Stable Diffusion 31.69 0.32
MFTF 32.26 0.32
Table 1: Evaluation metrics comparing LPIPS and CLIP-Score for images with and without layout control.
Figure 7: MFTF concurrently control the positions and arrangements of multiple objects within a scene, enables independent adjustments to the layout of each object while preserving overall structural coherence.
@@ 4.4.1 Layout Control and Semantic Editing
The model successfully demonstrates concurrent single-object positional control and semantic editing, as illustrated in Fig. 6. The first column presents images generated from the source prompts: “a cat sitting on a chair” and ”a painting on the wall in a cozy house”. Following the application of layout controls $L$, the resulting images are shown in the second, third. These results highlight the ability to apply positional control, including translation, rotation, and scaling, both independently and simultaneously for target image generation. In the fourth columns of Fig. 6, object-level layout control and semantic editing are applied together, with modified target prompts such as “a dog sitting on a chair” and “a painting above desk in a cozy house”.
@@ 4.4.2 Multiple Object Layout Control
The model is capable of controlling the positions of multiple objects simultaneously, as demonstrated in Fig. 7. The first image in the first row, generated from the source prompt $P_{s}$ “a small car and a cabin in the style of Disney, artstation”, shows the initial configuration. In this case, the “car” on the left-hand side is translated to the left, while the “cabin” is translated to the right. Subsequently, the “car” is rotated by $L=30^{\circ}$ and $L=330^{\circ}$ clockwise in the second to fourth images. The second row illustrates another example of concurrent object position control for multiple objects.
@@
4.5 Ablation Study
The ablation study investigated the effects of layout control across denoising steps and self-attention layers. As illustrated in Fig. 12, the early denoising steps play a critical role in determining object positioning and arrangement, while the later steps refine structural details. Effective layout control requires inclusion of the initial denoising steps; omission of these steps results in a failure to achieve the desired positional control. The layer control ablation study, shown in Fig. 13, highlights that applying the attention mask to higher-level layers achieves effective layout control. Conversely, when the mask is applied to fewer self-attention layers, the control effectiveness decreases significantly, ultimately leading to a complete loss of positional control. These findings emphasize the importance of both step and layer selection in achieving robust layout control.
4.6 Limitations
A primary challenge arises from the mixed structural information present in the cross-attention layer, which makes it difficult to fully encapsulate complete structural information within a single attention mask. This limitation hinders the model’s ability to fully separate objects from the background. Additionally, in multi-object position control, the model struggles with separate objects with complex interrelationships, particularly in scenarios involving tightly coupled or overlapping objects.
5 Conclusion
In this paper, we have introduced MFTF, a mask-free, fine-tuning-free diffusion model for object-level layout control. It enables both single-object and multi-object layout control, supports concurrent layout manipulation and semantic editing, and facilitates text-guided image segmentation. These capabilities showcase its promising potential for advancing research and applying it could empower users to custom images generations with unprecedented ease.
Supplementary Material
6 Supplementary
6.1 Modified multi-self-attention
Figure 8: Modified multi-self-attention.
6.2 Cross-attention Map
The cross-attention maps corresponding to individual tokens $i$ are illustrated in SFig. 9, which encapsulates both semantic and structural information. However, the structural information corresponding to each text token overlaps significantly, which complicates achieving precise absolute layout control. For instance, the cross-attention map in SFig. 9 generated for the prompt “a small cat sitting on a desk”, the token “cat” captures the majority of the structural information for the cat object. However, other tokens such as “small”, “sitting”, “on”, and “desk” also contribute overlapping profile information related to the “cat”, thereby demonstrating the inherent redundancy in the attention map.
6.3 Self-attention Query
Figure 10: Visualization of $Q$ is presented at different self-attention layers $l=5,10,15$, showing masks from the following conditions: without cross-attention mask, with cross-attention mask applied, and with additional positional control $L$.
6.4 Text-Controlled Semantic Segmentation
@@
The attention map $M_{s}$ and the masked queries, $Q_{s}$ or $Q_{\text{background}}$, can also be utilized for text-guided image segmentation, as illustrated in SFig. 11. Here, $M_{s}$ serves as the mask for the token ”panda”, which emphasizes the object’s profile, while $Q_{s}$ or $Q_{\text{background}}$ is used to generate detailed feature segmentation.
Figure 11: MFTF model enables text-controlled image segmentation through the generation of segmentation masks. Segmentation cross-attention mask $M_{s}$ and the masked query $Q_{ms}$ for token “panda” and “bird”.
6.5 Ablation Object-level Layout Control